# Improving Pareto Set Learning for Expensive Multi-objective Optimization via Stein Variational Hypernetworks

### v1

@@ Abstract
Expensive multi-objective optimization problems (EMOPs) are common in real-world scenarios where evaluating objective functions is costly and involves extensive computations or physical experiments. Current Pareto set learning methods for such problems often rely on surrogate models like Gaussian processes to approximate the objective functions. These surrogate models can become fragmented, resulting in numerous small uncertain regions between explored solutions. When using acquisition functions such as the Lower Confidence Bound (LCB), these uncertain regions can turn into pseudo-local optima, complicating the search for globally optimal solutions. To address these challenges, we propose a novel approach called SVH-PSL, which integrates Stein Variational Gradient Descent (SVGD) with Hypernetworks for efficient Pareto set learning. Our method addresses the issues of fragmented surrogate models and pseudo-local optima by collectively moving particles in a manner that smooths out the solution space. The particles interact with each other through a kernel function, which helps maintain diversity and encourages the exploration of underexplored regions. This kernel-based interaction prevents particles from clustering around pseudo-local optima and promotes convergence towards globally optimal solutions. Our approach aims to establish robust relationships between trade-off reference vectors and their corresponding true Pareto solutions, overcoming the limitations of existing methods. Through extensive experiments across both synthetic and real-world MOO benchmarks, we demonstrate that SVH-PSL [^1] Code is available at https://github.com/nguyenduc810/SVH-PSL significantly improves the quality of the learned Pareto set, offering a promising solution for expensive multi-objective optimization problems.
@@ Introduction
We sample $K$ preference vectors $\bm{r}_{i}$ and fit them to the Pareto set model $h(r|\theta)$ to generate solutions $\{\bm{x}_{r_{i}}\}_{i=1}^{K}$. For a problem with $\bm{n}$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\bm{x}_{r_{i}})=[f_{1}(\bm{x}_{r_{i}}),...,f_{n}(\bm{x}_{r_{i}})]$, where $\mathcal{F}_{i}$ is considered as a particle. The set of particles $\mathbf{F}={\{\mathcal{F}_{i}}\}^{K}_{i=1}$ serves as the initialization set. We iteratively move the particles towards the Pareto front using the update rule:
@@
$LHV(\mathcal{P}_{f},\hat{\mathcal{P}}_{f})=\log\Big{(}\text{HV}\big{(}\mathcal{P}_{f}\big{)}-\text{HV}\big{(}\hat{\mathcal{P}}_{f}\big{)}\Big{)}$
@@
Experiment Setting
For all experiments, we randomly generate 20 initial solutions for expensive evaluations. The model is trained over 20 iterations with a batch size of b = 5. In Formula 14, the hyperparameters are set as $c=med$ (the median of the pairwise distances between samples) and $\alpha=0.1$. Each problem is trained 5 times to eliminate randomness.
@@
Figure 4: Illustration of SVH-PSL performance with and without local kernel integration in real-world problems which have a complex PF.
@@
The trade-off between exploration and exploitation. In Formula 14, the second term acts as the repulsion component, with the parameter $\alpha$ serving as a trade-off between exploration and exploitation. As shown in Figure 5, $\alpha$ significantly impacts performance. When $\alpha$ is small (5(a)), the variance is high, indicating that insufficient repulsion makes it difficult for the particles to escape pseudo-optimal points, leading to instability. Conversely, a larger $\alpha$ (5(b)) encourages exploration, driving the particles toward the optimal solution.
(a) Small $\alpha$.
(b) Large $\alpha$.
(c) Approximate PF (Small $\alpha$).
(d) Approximate PF (Large $\alpha$).
Figure 5: Performance on ZDT1 with varying values of alpha.
@@
Figure 6: Performance Comparison in High-Dimensional ZDT1 and ZDT2 Problem

### v2

@@ Abstract
Expensive multi-objective optimization problems (EMOPs) are common in real-world scenarios where evaluating objective functions is costly and involves extensive computations or physical experiments. Current Pareto set learning methods for such problems often rely on surrogate models like Gaussian processes to approximate the objective functions. These surrogate models can become fragmented, resulting in numerous small uncertain regions between explored solutions. When using acquisition functions such as the Lower Confidence Bound (LCB), these uncertain regions can turn into pseudo-local optima, complicating the search for globally optimal solutions. To address these challenges, we propose a novel approach called SVH-PSL, which integrates Stein Variational Gradient Descent (SVGD) with Hypernetworks for efficient Pareto set learning. Our method addresses the issues of fragmented surrogate models and pseudo-local optima by collectively moving particles in a manner that smooths out the solution space. The particles interact with each other through a kernel function, which helps maintain diversity and encourages the exploration of underexplored regions. This kernel-based interaction prevents particles from clustering around pseudo-local optima and promotes convergence towards globally optimal solutions. Our approach aims to establish robust relationships between trade-off reference vectors and their corresponding true Pareto solutions, overcoming the limitations of existing methods. Through extensive experiments across both synthetic and real-world MOO benchmarks, we demonstrate that SVH-PSL significantly improves the quality of the learned Pareto set, offering a promising solution for expensive multi-objective optimization problems.
@@ Introduction
We sample $K$ preference vectors $\bm{r}_{i}$ and fit them to the Pareto set model $h(r|\theta)$ to generate solutions $\{\bm{x}_{r_{i}}\}_{i=1}^{K}$. For a problem with $\bm{m}$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\bm{x}_{r_{i}})=[f_{1}(\bm{x}_{r_{i}}),...,f_{m}(\bm{x}_{r_{i}})]$, where $\mathcal{F}_{i}$ is considered as a particle. The set of particles $\mathbf{F}={\{\mathcal{F}_{i}}\}^{K}_{i=1}$ serves as the initialization set. We iteratively move the particles towards the Pareto front using the update rule:
@@
$LHD(\mathcal{P}_{f},\hat{\mathcal{P}}_{f})=\log\Big{(}\text{HV}\big{(}\mathcal{P}_{f}\big{)}-\text{HV}\big{(}\hat{\mathcal{P}}_{f}\big{)}\Big{)}$
@@
Experiment Settings
For all experiments [^1] Code is available at https://github.com/nguyenduc810/SVH-PSL, we randomly generate 20 initial solutions for expensive evaluations. The model is trained over 20 iterations with a batch size of b = 5. In Formula 14, the hyperparameters are set as $c=med$ (the median of the pairwise distances between samples) and $\alpha=0.1$. Each problem is trained 5 times to eliminate randomness.
@@
Figure 4: Illustration of SVH-PSL performance with and without local kernel integration in real-world problems with a complex PF.
@@
The Trade-Off Between Exploration and Exploitation. In Formula 14, the second term acts as the repulsion component, with the parameter $\alpha$ serving as a trade-off between exploration and exploitation. As shown in Figure 5, $\alpha$ significantly impacts performance. When $\alpha$ is small (left), the variance is high, indicating that insufficient repulsion makes it difficult for the particles to escape pseudo-optimal points, leading to instability. Conversely, a larger $\alpha$ (right) encourages exploration, driving the particles toward the optimal solution.
@@
Figure 5: Impact of $\alpha$ on the trade-off between exploration and exploitation in ZDT1 (LHD & Pareto front).
Figure 6: Performance comparison of high-dimensional ZDT1 and ZDT2 problems.